OpenStudy (anonymous):

Implicit differentiation: ycosx=x^2+y^2 what am i doing wrong here? derivative of both sides: =((y)(-sinx)+(cosx)(dy/dx)=2x+2y(dy/dx) =(-ysinx+cosx(dy/dx)=2x+2y(dy/dx) =cosx(dy/dx) - 2y(dy/dx) = ysinx+2x =(cosx-2y)dy/dx = ysinx+2x =dy/dx = (ysinx+2x)/(cosx-2y)

OpenStudy (anonymous):

It looks right to me.

OpenStudy (anonymous):

nick....honey.....dear.....any other form of weird name for reference......i have absolutely no clue what language this question is in

OpenStudy (anonymous):

you just call me honey lol

OpenStudy (anonymous):

is that really all you picked up from what i posted lol

OpenStudy (anonymous):

Implicit differentiation: ycosx=x^2+y^2 ((y)(-sinx)+(cosx)(dy/dx)=2x+2y(dy/dx) (-ysinx+cosx(dy/dx)=2x+2y(dy/dx) collect terms cosx(dy/dx)-2y(dy/dx)=2x+ysinx <---------- (dy/dx)(cosx-2y)=2x+ysinx dy/dx))=(2x+ysinx)/(cosx-2y

OpenStudy (anonymous):

seems like you did it correctly...lol :D

OpenStudy (anonymous):

im starting to think my book might have some errors i appreciate mark and johnny

OpenStudy (anonymous):

YW....ok maybe its the book since you done it right ..lol have fun n good luck now